# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ 4.4.3 Sketches of proofs
The Gromov hyperbolic group $\Gamma_{0}$ has the following property: there exist a finite subset $S$ of $\Gamma$ and a constant $C^{\prime}>0$ such that for any $\gamma\in\Gamma_{0}$ we can find $s\in S$ with $\mathrm{transl}_{\mathrm{Cay}(\Gamma_{0})}(s\gamma)\geq\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)-C^{\prime}$. (If $\Gamma_{0}$ is nonelementary, then we can take $S=\{\gamma_{1}^{N},\gamma_{1}^{-N},\gamma_{2}^{N},\gamma_{2}^{-N}\}$ for some large $N$, where $\gamma_{1},\gamma_{2}\in\Gamma_{0}$ are infinite-order elements with pairwise distinct attracting and repelling fixed points in $\partial_{\infty}\Gamma_{0}$: see e.g. [137, Lem. B.2].)
@@
We have seen in the proof of (1$\Rightarrow$ (2) that for any $m\in\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, the orbital map $\gamma\mapsto\rho(\gamma)\cdot m$ is a quasi-isometry from $\Gamma_{0}$ to $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. It is a classical result in geometric group theory (see e.g. [64, Th. 11.108]) that such a quasi-isometry extends to a $\Gamma_{0}$-equivariant homeomorphism $\xi$ from $\partial_{\infty}\Gamma_{0}$ to $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. Here $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ is a subset of $\partial_{\infty}X$ (namely the intersection of $\partial_{\infty}X$ with the closure of $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ in $\overline{X}$). Thus we can view $\xi$ as a $\rho$-equivariant, continuous, injective map from $\partial_{\infty}\Gamma_{0}$ to $\partial_{\infty}X$, such that for any $(\gamma_{k})\in\Gamma_{0}^{\mathbb{N}}$ and any $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$, if $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$, then $\rho(\gamma_{k})\cdot z\to\xi(w^{+})$ for all $z\in\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\smallsetminus\{\xi(w^{-})\}$.
In order to see that this last convergence holds for all $z\in\partial_{\infty}X\smallsetminus\{\xi(w^{-})\}$, one possibility is to use the fact (Cartan decomposition) that if we choose a point $x\in X$ and a geodesic line $\mathcal{G}$ of $X$ through $x$ with endpoints $z^{+},z^{-}\in\partial_{\infty}X$, then any element $\rho(\gamma_{k})\in G$ can be written as $\rho(\gamma_{k})=\kappa_{k}a_{k}\kappa^{\prime}_{k}$ where $\kappa_{k},\kappa^{\prime}_{k}\in G$ fix $x$ and $a_{k}\in G$ is a pure translation along $\mathcal{G}$ towards $z^{+}$. The subgroup of $G$ fixing $x$ is compact (it is conjugate to $K$); therefore, up to passing to a subsequence we may assume that $(\kappa_{k})_{k\in\mathbb{N}},(\kappa^{\prime}_{k})_{k\in\mathbb{N}}$ converge respectively to some $\kappa,\kappa^{\prime}\in G$. Since $\rho$ has finite kernel and discrete image, we have $a_{k}\cdot z\to z^{+}$ for all $z\in\partial_{\infty}X\smallsetminus\{z^{-}\}$. Therefore $\rho(\gamma_{k})\cdot z\to\kappa\cdot z^{+}$ for all $z\in\partial_{\infty}X\smallsetminus\{{\kappa^{\prime}}^{-1}\cdot z^{-}\}$. Necessarily $\kappa\cdot z^{+}=\xi(w^{+})$ and ${\kappa^{\prime}}^{-1}\cdot z^{-}=\xi(w^{-})$. ∎
The implication (5$\Rightarrow$ (4) is immediate by taking $\gamma_{k}=\gamma^{k}$ for some infinite-order element $\gamma\in\Gamma_{0}$. The implication (4$\Rightarrow$ (2) can be proved using flows as in Section 5.1 below (see Remark 4.29 and the implication (1$\Rightarrow$ (2) in Theorem 4.37).
@@
Suppose by contradiction that the action of $\Gamma_{0}$ on $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ via $\rho$ is not cocompact. Let $(\varepsilon_{m})_{m\in\mathbb{N}}$ be a sequence of positive reals going to 00. For any $m$, the set $\mathcal{K}_{m}:=\{x\in\nolinebreak\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\,|\,\mathtt{d}_{\mathrm{Euc}}(x,\Lambda_{\Gamma})\geq\nolinebreak\varepsilon_{m}\}$ is compact, hence there exists a $\rho(\Gamma_{0})$-orbit contained in $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\smallsetminus\mathcal{K}_{m}$. By proper discontinuity of the action on $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, the supremum of $\mathtt{d}_{\mathrm{Euc}}(\cdot,\Lambda_{\Gamma})$ on this orbit is achieved at some point $x_{m}\in\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, and by construction we have $0<\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot x_{m},\Lambda_{\Gamma})\leq\mathtt{d}_{\mathrm{Euc}}(x_{m},\Lambda_{\Gamma})\leq\varepsilon_{m}$ for all $\gamma\in\Gamma_{0}$. Up to passing to a subsequence, we may assume that $(x_{m})_{m\in\mathbb{N}}$ converges to some $z\in\Lambda_{\Gamma}$. Consider a neighbourhood $\mathcal{U}$ of $z$ in $\mathbb{R}^{n}$ and an element $\gamma\in\Gamma_{0}$ such that (5) holds, and let $c>1$ be the infimum in (5). For any $m\in\mathbb{N}$, there exists $z_{m}\in\Lambda_{\Gamma}$ such that $\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\Gamma})=\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot x_{m},\rho(\gamma)\cdot z_{m})$. For large enough $m$ we have $x_{m},z_{m}\in\mathcal{U}$, and so $\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\Gamma})\geq c\,\mathtt{d}_{\mathrm{Euc}}(x_{m},z_{m})\geq c\,\mathtt{d}_{\mathrm{Euc}}(x_{m},\Lambda_{\Gamma})\geq c\,\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\Gamma})>0$. This is impossible since $c>1$. ∎
@@ 5 Classes of discrete subgroups in higher real rank
We have seen in Section 4 two important classes of discrete subgroups of semisimple Lie groups $G$ with $\mathrm{rank}_{\mathbb{R}}(G)=1$, namely convex cocompact subgroups and geometrically finite subgroups. These classes have been much studied, although many interesting questions remain even in the case of $G=\mathrm{PO}(n,1)$ for $n\geq 4$ (see e.g. [93]).
We now turn to infinite discrete subgroups of semisimple Lie groups $G$ for $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$. These discrete subgroups, beyond lattices, remain much more mysterious. Recently, an important class has emerged, namely the class of Anosov subgroups, which are by definition the images of the Anosov representations of Gromov hyperbolic groups introduced by Labourie [107] as part of his study of Hitchin representations (see Section 3.3). In fact, most examples in Section 3 are Anosov subgroups. We now discuss these subgroups, make the link with convex cocompactness, and mention some generalisations.
5.1 Anosov subgroups
@@
5.1.1 Definition and first observations
@@
  • satisfy a uniform contraction property (Condition 4.28 below) which strengthens the dynamics-preserving condition of Theorem 4.23.(4).
By the dynamics-preserving condition of Theorem 4.23.(4) for $\xi_{i}$ we mean that for any infinite-order element $\gamma\in\Gamma_{0}$, the image by $\xi_{i}$ of the attracting fixed point of $\gamma$ in $\partial_{\infty}\Gamma_{0}$ (see Section 4.4) is an attracting fixed point of $\rho(\gamma)$ in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
We note that for an element $g\in G$, the property of admitting an attracting fixed point in $G/P_{i}$ can be characterised in terms of eigenvalues, namely as $(\lambda_{i}-\lambda_{i+1})(g)>0$ (Notation 4.34). In this case the attracting fixed point is unique and we say that $g$ is proximal in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
@@
The uniform contraction property in Definition 4.25 is reminiscent of the condition defining Anosov flows in dynamics, which explains the terminology Anosov representation. Before stating it (Condition 4.28), let us make a few elementary observations that already follow from the fact that $\xi_{i}$ and $\xi_{d-i}$ are continuous, transverse, and dynamics-preserving.
@@
(1) Recall from Section 4.4 that the subset of $\partial_{\infty}\Gamma_{0}$ consisting of the attracting fixed points of infinite-order elements of $\Gamma_{0}$ is dense in $\partial_{\infty}\Gamma_{0}$. Since $\xi_{i}$ and $\xi_{d-i}$ are dynamics-preserving, they are uniquely determined on this subset, and compatible on this subset. By continuity, they are uniquely determined and compatible on all of $\partial_{\infty}\Gamma_{0}$. Moreover, the image of $\xi_{i}$ is the proximal limit set of $\rho(\Gamma_{0})$ in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
@@
(3) Suppose $\Gamma_{0}$ is nonelementary. In order to show that $\rho$ has finite kernel and discrete image, it is sufficient to consider an arbitrary sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct points of $\Gamma_{0}$ and to check that $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ does not converge to the identity of $G$. Recall from Section 4.4 that the action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$ is a convergence action. Therefore, up to passing to a subsequence, there exist $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$ such that $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$. By $\rho$-equivariance and continuity of $\xi_{i}$, we then have $\rho(\gamma_{k})\cdot\xi_{i}(w)=\xi_{i}(\gamma_{k}\cdot w)\to\xi_{i}(w^{+})$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$. Since $\partial_{\infty}\Gamma_{0}$ is infinite and $\xi_{i}$ is injective, there exists $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$ such that $\xi_{i}(w)\neq\xi_{i}(w^{+})$. The convergence $\rho(\gamma_{k})\cdot\xi_{i}(w)\to\xi_{i}(w^{+})$ then implies that $(\rho(\gamma_{k}))_{n\in\mathbb{N}}$ does not converge to the identity of $G$. This shows that $\rho$ has finite kernel and discrete image.
@@
5.1.2 The uniform contraction condition
@@
For simplicity, we take $G=\mathrm{SL}^{\pm}(d,\mathbb{K})$ (see Remark 4.26). Given a representation $\rho:\Gamma_{0}\to G$, let
@@
be the natural flat vector bundle over $T^{1}(M)=\Gamma_{0}\backslash T^{1}(\widetilde{M})$ associated to $\rho$, where $\Gamma_{0}$ acts on $T^{1}(\widetilde{M})\times\mathbb{K}^{d}$ by $\gamma\cdot(\tilde{x},v)=(\gamma\cdot\tilde{x},\rho(\gamma)\cdot v)$. The geodesic flow $(\varphi_{t})_{t\in\mathbb{R}}$ on $T^{1}(M)$ lifts to a flow $(\psi_{t})_{t\in\mathbb{R}}$ on $E^{\rho}$, given by $\psi_{t}\cdot[(\tilde{x},v)]=[(\varphi_{t}\cdot\tilde{x},v)]$.
For each $\tilde{x}\in\nolinebreak T^{1}(\widetilde{M})$, the transversality of the boundary maps induces a decomposition $\mathbb{K}^{d}=\xi_{i}(\tilde{x}^{+})\oplus\xi_{d-i}(\tilde{x}^{-})$, where $\tilde{x}^{\pm}=\lim_{t\to\pm\infty}\varphi_{t}\cdot\tilde{x}$ are the forward and backward endpoints of the geodesic determined by $\tilde{x}$, and this defines a decomposition of the vector bundle $E^{\rho}$ into the direct sum of two subbundles $E_{i}^{\rho}=\{[(\tilde{x},v)]\,|\,v\in\xi_{i}(\tilde{x}^{+})\}$ and $E_{d-i}^{\rho}=\{[(\tilde{x},v)]\,|\,v\in\nolinebreak\xi_{d-i}(\tilde{x}^{-})\}$. This decomposition is invariant under the flow $(\psi_{t})$. By definition, the representation $\rho$ is $P_{i}$-Anosov if the following “dominated splitting” condition is satisfied.
Condition 4.28.
The flow $(\psi_{t})_{t\in\mathbb{R}}$ uniformly contracts $E_{i}^{\rho}$ with respect to $E_{d-i}^{\rho}$, i.e. there exist $C,C^{\prime}>0$ such that for any $t\geq 0$, any $x\in T^{1}(M)$, and any nonzero $\upsilon_{i}\in E_{i}^{\rho}(x)$ and $\upsilon_{d-i}\in E_{d-i}^{\rho}(x)$,
@@
where $(\|\cdot\|_{x})_{x\in T^{1}(M)}$ is any fixed continuous family of norms on the fibers $E^{\rho}(x)$.
@@
Guichard and Wienhard [82] showed that if there exist $\rho$-equivariant maps $\xi_{i}$ and $\xi_{d-i}$ which are continuous, transverse, and dynamics-preserving, and if the group $\rho(\Gamma_{0})$ is Zariski-dense in $G$, then condition 4.28 is automatically satisfied.
5.1.3 Properties
  • $P_{i}$-Anosov is equivalent to $P_{d-i}$-Anosov, as the integers $i$ and $d-i$ play a similar role in Definition 4.25 and Condition 4.28 (up to reversing the flow, which switches contraction and expansion). In particular, we may restrict to $P_{i}$-Anosov for $21\leq i\leq d/$ 2.
@@ 5.1.3 Properties
  • When $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$ (i.e. $d\geq 3$ for $G=\mathrm{PGL}(d,\mathbb{K})$ or $\mathrm{SL}^{\pm}(d,\mathbb{K})$), Anosov subgroups are not lattices of $G$ (since Anosov subgroups are Gromov hyperbolic unlike lattices, see Section 2.3).
  • Uniform contraction over a compact space as in Condition 4.28 is stable under small deformations, which implies the following analogue of Fact 4.18.
    Fact 4.30.
@@
  • Anosov representations behave well under Lie group homomorphisms: the following holds similarly to Fact 4.19. (We refer to Remark 4.24 for the notion of a hyperbolic element of $G^{\prime}$.)
    Fact 4.31 (see [82]).

### 28 Mar 2024 v2

@@ 4 A few characterisations of convex cocompactness
The Gromov hyperbolic group $\Gamma_{0}$ has the following property: there exist a finite subset $S$ of $\Gamma_{0}$ and a constant $C^{\prime}>0$ such that for any $\gamma\in\Gamma_{0}$ we can find $s\in S$ with $\mathrm{transl}_{\mathrm{Cay}(\Gamma_{0})}(s\gamma)\geq\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)-C^{\prime}$. (If $\Gamma_{0}$ is nonelementary, then we can take $S=\{\gamma_{1}^{N},\gamma_{1}^{-N},\gamma_{2}^{N},\gamma_{2}^{-N}\}$ for some large $N$, where $\gamma_{1},\gamma_{2}\in\Gamma_{0}$ are infinite-order elements such that the attracting fixed points in $\partial_{\infty}\Gamma_{0}$ of $\gamma_{1}$, $\gamma_{1}^{-1}$, $\gamma_{2}$, and $\gamma_{2}^{-1}$ are pairwise distinct: see e.g. [139, Lem. B.2].)
@@
We have seen in the proof of (1$\Rightarrow$ (2) that for any $m\in\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, the orbital map $\gamma\mapsto\rho(\gamma)\cdot m$ is a quasi-isometry from $\Gamma_{0}$ to $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. It is a classical result in geometric group theory (see e.g. [65, Th. 11.108]) that such a quasi-isometry extends to a $\Gamma_{0}$-equivariant homeomorphism $\xi$ from $\partial_{\infty}\Gamma_{0}$ to $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$. Here $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ is a subset of $\partial_{\infty}X$ (namely the intersection of $\partial_{\infty}X$ with the closure of $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ in $\overline{X}$). Thus we can view $\xi$ as a $\rho$-equivariant, continuous, injective map from $\partial_{\infty}\Gamma_{0}$ to $\partial_{\infty}X$, such that for any $(\gamma_{k})\in\Gamma_{0}^{\mathbb{N}}$ and any $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$, if $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$, then $\rho(\gamma_{k})\cdot z\to\xi(w^{+})$ for all $z\in\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\smallsetminus\{\xi(w^{-})\}$.
In order to see that this last convergence holds for all $z\in\partial_{\infty}X\smallsetminus\{\xi(w^{-})\}$, one possibility is to use the fact (Cartan decomposition) that if we choose a point $x\in X$ and a geodesic line $\mathcal{G}$ of $X$ through $x$ with endpoints $z_{0}^{+},z_{0}^{-}\in\partial_{\infty}X$, then any element $\rho(\gamma_{k})\in G$ can be written as $\rho(\gamma_{k})=\kappa_{k}a_{k}\kappa^{\prime}_{k}$ where $\kappa_{k},\kappa^{\prime}_{k}\in G$ fix $x$ and $a_{k}\in G$ is a pure translation along $\mathcal{G}$ towards $z_{0}^{+}$. The subgroup of $G$ fixing $x$ is compact (it is conjugate to $K$); therefore, up to passing to a subsequence we may assume that $(\kappa_{k})_{k\in\mathbb{N}},(\kappa^{\prime}_{k})_{k\in\mathbb{N}}$ converge respectively to some $\kappa,\kappa^{\prime}\in G$. Since $\rho$ has finite kernel and discrete image, we have $a_{k}\cdot z\to z_{0}^{+}$ for all $z\in\partial_{\infty}X\smallsetminus\{z_{0}^{-}\}$. Therefore $\rho(\gamma_{k})\cdot z\to\kappa\cdot z_{0}^{+}$ for all $z\in\partial_{\infty}X\smallsetminus\{{\kappa^{\prime}}^{-1}\cdot z_{0}^{-}\}$. Necessarily $\kappa\cdot z_{0}^{+}=\xi(w^{+})$ and ${\kappa^{\prime}}^{-1}\cdot z_{0}^{-}=\xi(w^{-})$. ∎
The implication (5$\Rightarrow$ (4) is immediate by considering, for any infinite-order element $\gamma\in\Gamma_{0}$, the sequence $(\gamma_{k}):=(\gamma^{k})\in\Gamma_{0}^{\mathbb{N}}$. The implication (4$\Rightarrow$ (2) can be proved using flows as in Section 1 below (see Remark 5.5 and the implication (1$\Rightarrow$ (2) in Theorem 5.13).
@@
Suppose by contradiction that the action of $\Gamma_{0}$ on $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ via $\rho$ is not cocompact. Let $(\varepsilon_{m})_{m\in\mathbb{N}}$ be a sequence of positive reals going to 00. For any $m$, the set $\mathcal{K}_{m}:=\{x\in\nolinebreak\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\,|\,\mathtt{d}_{\mathrm{Euc}}(x,\Lambda_{\rho(\Gamma_{0})})\geq\nolinebreak\varepsilon_{m}\}$ is compact, hence there exists a $\rho(\Gamma_{0})$-orbit contained in $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\smallsetminus\mathcal{K}_{m}$. By proper discontinuity of the action on $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, the supremum of $\mathtt{d}_{\mathrm{Euc}}(\cdot,\Lambda_{\rho(\Gamma_{0})})$ on this orbit is achieved at some point $x_{m}\in\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, and by construction we have $0<\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot x_{m},\Lambda_{\rho(\Gamma_{0})})\leq\mathtt{d}_{\mathrm{Euc}}(x_{m},\Lambda_{\rho(\Gamma_{0})})\leq\varepsilon_{m}$ for all $\gamma\in\Gamma_{0}$. Up to passing to a subsequence, we may assume that $(x_{m})_{m\in\mathbb{N}}$ converges to some $z\in\Lambda_{\rho(\Gamma_{0})}$. Consider a neighbourhood $\mathcal{U}$ of $z$ in $\mathbb{R}^{n}$ and an element $\gamma\in\Gamma_{0}$ such that (5) holds for $\mathtt{d}_{\mathrm{Euc}}$, and let $c>1$ be the infimum in (5). For any $m\in\mathbb{N}$, there exists $z_{m}\in\Lambda_{\rho(\Gamma_{0})}$ such that $\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\rho(\Gamma_{0})})=\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot x_{m},\rho(\gamma)\cdot z_{m})$. For large enough $m$ we have $x_{m},z_{m}\in\mathcal{U}$, and so $\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\rho(\Gamma_{0})})\geq c\,\mathtt{d}_{\mathrm{Euc}}(x_{m},z_{m})\geq c\,\mathtt{d}_{\mathrm{Euc}}(x_{m},\Lambda_{\rho(\Gamma_{0})})\geq c\,\mathtt{d}_{\mathrm{Euc}}(\rho(\gamma)\cdot\nolinebreak x_{m},\Lambda_{\rho(\Gamma_{0})})>0$. This is impossible since $c>1$. ∎
@@ 5 Classes of discrete subgroups in higher real rank
We have seen in Section 4 two important classes of discrete subgroups of semisimple Lie groups $G$ with $\mathrm{rank}_{\mathbb{R}}(G)=1$, namely convex cocompact subgroups and geometrically finite subgroups. These classes have been much studied, although many interesting questions remain open even in the case of $G=\mathrm{PO}(n,1)$ for $n\geq 4$ (see e.g. [95]).
We now turn to infinite discrete subgroups of semisimple Lie groups $G$ for $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$. These discrete subgroups, beyond lattices, remain more mysterious, and very few general results are known (see [71] for a notable exception). Recently, an important class has emerged, namely the class of Anosov subgroups, which are by definition the images of the Anosov representations of Gromov hyperbolic groups introduced by Labourie [109] as part of his study of Hitchin representations (see Section 3). In fact, most examples in Section 3 are Anosov subgroups. We now discuss these subgroups, make the link with convex cocompactness, and mention some generalisations.
1 Anosov subgroups
@@
Definition and first observations
@@
  • satisfy a uniform contraction property (Condition 5.4 below) which strengthens the dynamics-preserving condition of Theorem 4.12.(4).
By the dynamics-preserving condition of Theorem 4.12.(4) for $\xi_{i}$ we mean that for any infinite-order element $\gamma\in\Gamma_{0}$, the image by $\xi_{i}$ of the attracting fixed point of $\gamma$ in $\partial_{\infty}\Gamma_{0}$ (see Section 4) is an attracting fixed point of $\rho(\gamma)$ in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
We note that for an element $g\in G$, the property of admitting an attracting fixed point in $G/P_{i}$ can be characterised in terms of eigenvalues, namely as $(\lambda_{i}-\lambda_{i+1})(g)>0$ (Notation 5.10). In this case the attracting fixed point is unique and we say that $g$ is proximal in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
@@
The uniform contraction property in Definition 5.1 is reminiscent of the condition defining Anosov flows in dynamics, which explains the terminology Anosov representation. Before stating it (Condition 5.4), let us make a few elementary observations that already follow from the fact that $\xi_{i}$ and $\xi_{d-i}$ are continuous, transverse, and dynamics-preserving.
@@
(1) Recall from Section 4 that the subset of $\partial_{\infty}\Gamma_{0}$ consisting of the attracting fixed points of infinite-order elements of $\Gamma_{0}$ is dense in $\partial_{\infty}\Gamma_{0}$. Since $\xi_{i}$ and $\xi_{d-i}$ are dynamics-preserving, they are uniquely determined on this subset, and compatible on this subset. By continuity, they are uniquely determined and compatible on all of $\partial_{\infty}\Gamma_{0}$. Moreover, the image of $\xi_{i}$ is the proximal limit set of $\rho(\Gamma_{0})$ in $\mathrm{Gr}_{i}(\mathbb{K}^{d})$.
@@
(3) Suppose $\Gamma_{0}$ is nonelementary. In order to show that $\rho$ has finite kernel and discrete image, it is sufficient to consider an arbitrary sequence $(\gamma_{k})_{k\in\mathbb{N}}$ of pairwise distinct points of $\Gamma_{0}$ and to check that $(\rho(\gamma_{k}))_{k\in\mathbb{N}}$ does not converge to the identity of $G$. Recall from Section 4 that the action of $\Gamma_{0}$ on $\partial_{\infty}\Gamma_{0}$ is a convergence action. Therefore, up to passing to a subsequence, there exist $w^{+},w^{-}\in\partial_{\infty}\Gamma_{0}$ such that $\gamma_{k}\cdot w\to w^{+}$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$. By $\rho$-equivariance and continuity of $\xi_{i}$, we then have $\rho(\gamma_{k})\cdot\xi_{i}(w)=\xi_{i}(\gamma_{k}\cdot w)\to\xi_{i}(w^{+})$ for all $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$. Since $\partial_{\infty}\Gamma_{0}$ is infinite and $\xi_{i}$ is injective, there exists $w\in\partial_{\infty}\Gamma_{0}\smallsetminus\{w^{-}\}$ such that $\xi_{i}(w)\neq\xi_{i}(w^{+})$. The convergence $\rho(\gamma_{k})\cdot\xi_{i}(w)\to\xi_{i}(w^{+})$ then implies that $(\rho(\gamma_{k}))_{n\in\mathbb{N}}$ does not converge to the identity element of $G$. This shows that $\rho$ has finite kernel and discrete image.
@@
The uniform contraction condition
@@
For simplicity, we take $G=\mathrm{SL}^{\pm}(d,\mathbb{K})$ (see Remark 5.2). Any representation $\rho:\Gamma_{0}\to G$ then determines a flat vector bundle
@@
over $T^{1}(M)=\Gamma_{0}\backslash T^{1}(\widetilde{M})$, where $\Gamma_{0}$ acts on $T^{1}(\widetilde{M})\times\mathbb{K}^{d}$ by $\gamma\cdot(\tilde{x},v)=(\gamma\cdot\tilde{x},\rho(\gamma)\cdot v)$. The geodesic flow $(\varphi_{t})_{t\in\mathbb{R}}$ on $T^{1}(M)$ lifts to a flow $(\psi_{t})_{t\in\mathbb{R}}$ on $E^{\rho}$, given by $\psi_{t}\cdot[(\tilde{x},v)]=[(\varphi_{t}\cdot\tilde{x},v)]$.
Suppose, as in Definition 5.1, that there exist continuous, transverse, $\rho$-equivariant boundary maps $\xi_{i}:\partial_{\infty}\Gamma_{0}\to\mathrm{Gr}_{i}(\mathbb{K}^{d})$ and $\xi_{d-i}:\partial_{\infty}\Gamma_{0}\to\mathrm{Gr}_{d-i}(\mathbb{K}^{d})$. By transversality, for each $\tilde{x}\in T^{1}(\widetilde{M})$ we have a decomposition $\mathbb{K}^{d}=\xi_{i}(\tilde{x}^{+})\oplus\xi_{d-i}(\tilde{x}^{-})$, where $\tilde{x}^{\pm}=\lim_{t\to\pm\infty}\varphi_{t}\cdot\tilde{x}\in\partial_{\infty}\widetilde{M}\simeq\partial_{\infty}\Gamma_{0}$ are the forward and backward endpoints of the geodesic determined by $\tilde{x}$, and this defines a decomposition of the vector bundle $E^{\rho}$ into the direct sum of two subbundles $E_{i}^{\rho}=\{[(\tilde{x},v)]\,|\,v\in\xi_{i}(\tilde{x}^{+})\}$ and $E_{d-i}^{\rho}=\{[(\tilde{x},v)]\,|\,v\in\nolinebreak\xi_{d-i}(\tilde{x}^{-})\}$. This decomposition is invariant under the flow $(\psi_{t})$. By definition, the representation $\rho$ is $P_{i}$-Anosov if the following “dominated splitting” condition is satisfied.
Condition 5.4.
The flow $(\psi_{t})_{t\in\mathbb{R}}$ uniformly contracts $E_{i}^{\rho}$ with respect to $E_{d-i}^{\rho}$, i.e. given a continuous family $(\|\cdot\|_{x})_{x\in T^{1}(M)}$ of norms on the fibers $E^{\rho}(x)$, there exist $C,C^{\prime}>0$ such that for any $t\geq 0$, any $x\in T^{1}(M)$, and any nonzero $\upsilon_{i}\in E_{i}^{\rho}(x)$ and $\upsilon_{d-i}\in E_{d-i}^{\rho}(x)$,
@@
Guichard and Wienhard [84] showed that if there exist $\rho$-equivariant maps $\xi_{i}$ and $\xi_{d-i}$ which are continuous, transverse, and dynamics-preserving, and if the group $\rho(\Gamma_{0})$ is Zariski-dense in $G$, then Condition 5.4 is automatically satisfied.
Properties
  • $P_{i}$-Anosov is equivalent to $P_{d-i}$-Anosov, as the integers $i$ and $d-i$ play a similar role in Definition 5.1 and Condition 5.4 (up to reversing the flow, which switches contraction and expansion). In particular, we may restrict to $P_{i}$-Anosov for $21\leq i\leq d/$ 2.
@@
  • When $\mathrm{rank}_{\mathbb{R}}(G)\geq 2$ (i.e. $d\geq 3$ for $G=\mathrm{PGL}(d,\mathbb{K})$ or $\mathrm{SL}^{\pm}(d,\mathbb{K})$), Anosov subgroups are not lattices of $G$ (since Anosov subgroups are Gromov hyperbolic unlike lattices, see Section 3).
  • Uniform contraction over a compact space as in Condition 5.4 is stable under small deformations, which implies the following analogue of Fact 4.7.
    Fact 5.6.
@@
  • Anosov representations behave well under Lie group homomorphisms: the following holds similarly to Fact 4.8. (We refer to Remark 4.13 for the notion of a hyperbolic element of $G^{\prime}$.)
    Fact 5.7 (see [84]).